How much pure alcohol should be added to a 40% alcohol solution to make 600mL of a 75%alcohol solution? (equations please)
Found 2 solutions by fcabanski, Alan3354:
Answer by fcabanski(1391)

We have to know how much there is of the 40% solution.

Answer by Alan3354(69443)

How much pure alcohol should be added to a 40% alcohol solution to make 600mL of a 75%alcohol solution?
--------------
p = pure
f = 40%
-----
100p + 40f = 600*75 (total solution)
p + f = 600 (total alcohol)
f = 600 - p
---
100p + 40*(600-p) = 45000
10p + 2400 - 4p = 4500
6p = 2100
p = 350 ml
f = 250 ml